OpenStudy (christos):

Is this correct, and if not where is the mistake? 21x^2 - 46x + 15 Factors of 21: 3 x 7 Factors of 15: 3 x 5 3 x 3 + 7 x 5 = 9 + 35 = 46 Since it has to be -46 and + 15, then the factors of 15 are both negative. Factored: (7x - 3)(3x - 5)

OpenStudy (christos):

Medal given!

OpenStudy (zehanz):

It gives after expanding 21x²-44x+15 :(

OpenStudy (zehanz):

Look at the mess WolframAlpha makes of it: http://www.wolframalpha.com/input/?i=factor+21x^2+-+46x+%2B+15 Seems there is an error in the thing.

OpenStudy (christos):

There is a mistake in the expression itself? So 21x^2 - 46x + 15 can't be factored?

OpenStudy (zehanz):

That's what I tried to say!

OpenStudy (anonymous):

9+35=44 not 46

OpenStudy (christos):

Nice one ingenuus, I didn't notice it. Do you have any solution for that?

OpenStudy (zehanz):

There is no solution for it. If the expression was 21x² - 44x + 15, everything would be fine.

OpenStudy (christos):

So isn't there another way/method for it to be factored?

OpenStudy (zehanz):

Take a look at the WolframAlpha solution...

OpenStudy (anonymous):

x-(23+\[21(x -(23+\sqrt{214})/21)(x-(23-\sqrt{214})/21)\])

OpenStudy (anonymous):

only the midde one is the answer..

OpenStudy (anonymous):

you couldn't have guessed that one could you :)

OpenStudy (anonymous):

the test shows it has 2 solutions, so it can be factorized. use the formula to figure it out.

OpenStudy (anonymous):

unfortunatly some irrational number will pop out

OpenStudy (christos):

So if I expand 21(x−(23+214−−−√)/21)(x−(23−214−−−√)/21) it's going to give me: 21x^2 - 46x + 15 ?

OpenStudy (christos):

or?

OpenStudy (anonymous):

why don't you try out yourself. i think that will be more helpful to you

OpenStudy (anonymous):

i just tried it and it works like magic :)

OpenStudy (anonymous):

you know, every ax^2+bx+c can be factorized if we allow complex numbers.

OpenStudy (anonymous):

anyways, i think you got my point right on the method of factorizing. good luck!

OpenStudy (christos):

So I will try to expand 21(x−(23+214−−−√)/21)(x−(23−214−−−√)/21) and try to get my result right?

OpenStudy (anonymous):

yeah it will get you right you know \[(a+b)(a-b)=a^2-b^2\] right? it will help you expand

OpenStudy (christos):

So with my way it can't be solved I take it

OpenStudy (anonymous):